# Pontryagin Neural Operator for Solving General-Sum Differential Games with Parametric State Constraints

### v1

@@ 5 Experiments and Results
We use the neural operators for closed-loop control on test cases uniformly sampled in $\mathcal{X}_{GT}$ and report their safety performance in percentage of collisions, where collisions are defined by $\boldsymbol{\theta}$. For better transparency, the comparison uses two sets of ground truth trajectories: The first contains 600 trajectories for each parameter configuration in $\Theta^{2}$; the second uses the same setting but only contains trajectories without physical collisions (for initial states from which collision cannot be avoided according to $\boldsymbol{\theta}=(1,1)$). Tab. 1 and Tab. 2 summarize safety performance comparisons with and without inevitable collisions, generalizing to testing cases across $\Theta^{2}$. The results show that PNO outperforms the hybrid neural operator in most cases. More importantly, it achieves this without relying on domain knowledge (informative trajectories in $\mathcal{X}_{GT}$). Trajectories from test player types are visualized in Fig. 2. It should be noted that in some cases neural operators perform better than the ground truth. This is because a multi-start BVP solver solves the ground truth trajectories with heuristic initial guesses. We also note that while BVP solutions are open-loop, their corresponding values are consistent with the HJI equations for this case study (see comparison between BVP and HJI value contours at $t=0$ in Fig. 4).
@@
Since branch and trunk nets learn a decomposition of the value function in the parameter space, Fig. 3 investigates the learned value structure through visualization of (1) the major basis value functions associated with the largest sample means of coefficients ($b_{k}$ in Eq. (5)) over $\Theta^{2}$ (Fig. 3a), (2) the ground truth values solved using level set for $\boldsymbol{\theta}=(1,1),~{}(1,5),~{}(5,1),~{}(5,5)$ (Fig. 3b), and (3) the sorted mean and standard deviation of absolute coefficients (Fig. 3c). We see that the major bases the mean and major variances of the value over $\Theta^{2}$ well. Further regularization on the branch net to promote basis sparsity and extension of existing generalization results to neural operators are worth investigating in future studies.
@@
Due to the reported importance of activation choice in PINN, we report ablation studies to understand the influence of activation on safety performance. Tab. 3 summarize the safety performance using different activation functions for comparisons w/ and w/o inevitable collisions. The comparison of closed-loop trajectories is visualized in Fig. 5. The results show that the choice of activation significantly affects the model performance: tanh outperforms sine and relu.
Figure 4: (a, b) Value contour of BVP solutions and HJI equations projected on $(d_{1},d_{2})$ with fixed $v_{1,2}=18m/s$ and $t=0$ for given $\boldsymbol{\theta}=(1,1)$, respectively. (c) Difference $|\vartheta_{BVP}-\vartheta_{HJI}|$ is visualized for $d_{1}$-$d_{2}$.
Figure 5: Closed-loop trajectories generated using PNO with (a) relu, (b) sine and (c) tanh activation functions without inevitable collisions.
Table 3: Safety performance comparison among different activation functions using PNO without inevitable collisions for each parameter configuration in $\Theta^{2}$
Player Types Activation Function Metrics relu  | sine  | tanh (1, 1) (1, 2) (1, 3) (1, 4) (1, 5) 81.3% 1.83% 0.33% 84.8% 4.33% 4.33% 86.2% 12.5% 5.67% 86.8% 5.33% 0.00% 87.8% 3.50% 1.00% (2, 1) (2, 2) (2, 3) (2, 4) (2, 5) 84.8% 4.33% 4.33% 89.2% 15.5% 2.83% 88.8% 12.0% 3.83% 90.3% 4.33% 1.17% 89.8% 0.67% 0.00% ($\theta_{1}$, $\theta_{2}$) (3, 1) (3, 2) (3, 3) (3, 4) (3, 5) Col.$\downarrow$ 86.2% 12.5% 5.67% 88.8% 12.0% 3.83% 92.5% 4.17% 0.67% 91.5% 0.83% 0.50% 93.7% 0.67% 0.67% (4, 1) (4, 2) (4, 3) (4, 4) (4, 5) 86.8% 5.33% 0.00% 90.3% 4.33% 1.17% 91.5% 0.83% 0.50% 94.8% 0.33% 0.00% 95.3% 0.17% 0.33% (5, 1) (5, 2) (5, 3) (5, 4) (5, 5) 87.8% 3.50% 1.00% 89.8% 0.67% 0.00% 93.7% 0.67% 0.67% 95.3% 0.17% 0.33% 95.8% 0.00% 0.00%
@@ 6 Conclusion
We introduced PNO to approximate parametric value functions with large Lipschitz constants as solutions to two-player general-sum differential games. Compared with the existing solution that requires informative supervisory data, PNO is fully self-supervised and free of prior knowledge about corner cases. For this reason, PNO is also free of convergence issues related to the existence of multiple equilibria or singular arcs in solving BVPs. In a case study involving two vehicles interacting at an uncontrolled intersection, our numerical results demonstrated that PNO yields superior safety performance compared to the hybrid neural operator, achieving this with a lower computational cost. The key ingredients of PNO are the self-supervised costate losses that align the learned values and value gradients with the discontinuous value dynamics governed by PMP, and the evolutionary sampling strategy that reveals informative state-space regions during the learning.

### v2

@@ 5 Experiments and Results
We use the neural operators to compute $\nabla_{\textbf{x}_{i}}\hat{\vartheta}_{i},\forall i=1,2$ as closed-loop control on test cases uniformly sampled in $\mathcal{X}_{GT}$ and report their safety performance in percentage of collisions, where collisions are defined by $\boldsymbol{\theta}$. For better transparency, the comparison uses two sets of ground truth trajectories: The first contains 600 trajectories for each parameter configuration in $\Theta^{2}$; the second uses the same setting but only contains trajectories without physical collisions (for initial states from which collision cannot be avoided according to $\boldsymbol{\theta}=(1,1)$). Tab. 1 and Tab. 2 summarize safety performance comparisons with and without inevitable collisions, generalizing to testing cases across $\Theta^{2}$. The results show that PNO outperforms the hybrid neural operator in most cases. More importantly, it achieves this without relying on domain knowledge (informative trajectories in $\mathcal{X}_{GT}$). Trajectories from test player types are visualized in Fig. 2. It should be noted that in some cases neural operators perform better than the ground truth. This is because a multi-start BVP solver solves the ground truth trajectories with heuristic initial guesses. We also note that while BVP solutions are open-loop, their corresponding values are consistent with the HJI equations for this case study (see comparison between BVP and HJI value contours at $t=0$ in Fig. 4).
@@
Since branch and trunk nets learn a decomposition of the value function in the parameter space, Fig. 3 investigates the learned value structure through visualization of (1) the major basis value functions associated with the 6 leading means of absolute coefficients ($b_{k}$ in Eq. (5)) over $\Theta^{2}$ (Fig. 3a), (2) the ground truth values solved using level set for $\boldsymbol{\theta}=(1,1),~{}(1,5),~{}(5,1),~{}(5,5)$ (Fig. 3b), and (3) the sorted mean and standard deviation of absolute coefficients over $\Theta^{2}$ along the output dimension of branch net, where the dominant basis value functions correspond to larger $|b_{k}|$ (Fig. 3c). Further regularization on the branch net to promote basis sparsity and extension of existing generalization results to neural operators are worth investigating in future studies.
@@
Due to the reported importance of activation choice in PINN, we conduct ablation studies to understand the influence of activation on safety performance. Tab. 3 summarizes the safety performance using different activation functions for comparisons w/ and w/o inevitable collisions. The comparison of closed-loop trajectories is visualized in Fig. 5. The results confirm that the choice of activation affects safety performance: tanh outperforms sin and relu.
Figure 4: (a, b) Value contour of BVP solutions and HJI equations projected respectively to $(d_{1},d_{2})$ with fixed $v_{1,2}=18m/s$ and $t=0$ for $\boldsymbol{\theta}=(1,1)$. (c) Difference $|\vartheta_{BVP}-\vartheta_{HJI}|$.
Figure 5: Closed-loop trajectories generated using PNO with (a) relu, (b) sin and (c) tanh activation functions without inevitable collisions.
Table 3: Safety comparison among activation functions using PNO w/o inevitable collisions for all configurations in $\Theta^{2}$
Player Types Activation Function Metrics relu  | sin  | tanh (1, 1) (1, 2) (1, 3) (1, 4) (1, 5) 81.3% 1.83% 0.17% 84.8% 4.33% 4.33% 86.2% 12.5% 5.67% 86.8% 5.33% 0.00% 87.8% 3.50% 1.00% (2, 1) (2, 2) (2, 3) (2, 4) (2, 5) 84.8% 4.33% 4.33% 88.0% 15.5% 2.83% 88.8% 12.0% 3.83% 90.3% 4.33% 1.17% 89.8% 0.67% 0.00% ($\theta_{1}$, $\theta_{2}$) (3, 1) (3, 2) (3, 3) (3, 4) (3, 5) Col.$\downarrow$ 86.2% 12.5% 5.67% 88.8% 12.0% 3.83% 92.5% 4.17% 0.67% 91.5% 0.83% 0.50% 93.7% 0.67% 0.67% (4, 1) (4, 2) (4, 3) (4, 4) (4, 5) 86.8% 5.33% 0.00% 90.3% 4.33% 1.17% 91.5% 0.83% 0.50% 94.8% 0.33% 0.00% 95.3% 0.17% 0.33% (5, 1) (5, 2) (5, 3) (5, 4) (5, 5) 87.8% 3.50% 1.00% 89.8% 0.67% 0.00% 93.7% 0.67% 0.67% 95.3% 0.17% 0.33% 95.8% 0.00% 0.00%
@@ 6 Conclusion
[…]